OpenStudy (anonymous):

evaluate cube root of -125 show steps / solution

OpenStudy (anonymous):

\[\sqrt[3]{-125}\]

OpenStudy (anonymous):

I'll give you a hint. \(-125 = -5^3\)

OpenStudy (anonymous):

\[\sqrt[3]{-125}=\sqrt[3]{-1}*\sqrt[3]{125}\]

OpenStudy (anonymous):

\[i*\sqrt[3]{125}= 5i\]

OpenStudy (anonymous):

5i = -5

OpenStudy (anonymous):

Uh.. The cubed root of -1 is not i.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You're thinking of square roots. You can take the cubed root of a negative number.

OpenStudy (anonymous):

\[\sqrt[3]{a^3} = a\]

OpenStudy (anonymous):

Therefore: \[\sqrt[3]{-125} = \sqrt[3]{-5^3} = ?\]

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

Indeed.

OpenStudy (anonymous):

so which is the correct solution then ?

OpenStudy (anonymous):

\[5\sqrt[3]{-1?}\]

OpenStudy (anonymous):

\[\sqrt[3]{-1}*\sqrt[3]{125}\]\[=5\sqrt[3]{-1}\]

OpenStudy (anonymous):

The cubed root of -1 is -1. It's just -5 for the answer.

OpenStudy (anonymous):

-1*5=-5